The data for the promotion status and academic qualification regarding 100 employees of a company is as follows:


Promotional. MBA. Academic qualification. Total

Status. Non - MBA


Promoted. 12. 48. 60

Not promoted. 18. 22. 40

Total. 30. 70. 100

At random one employee is picked up. What is the probability that:

1. He is MBA

2. He is promoted

3. He is promoted given that he is an MBA, and

4. He is an MBA given that he is promoted

5. He is promoted and MBA too?


1
Expert's answer
2021-05-07T08:57:03-0400
"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c}
 & MBA & Academic\\
& & qualifications & Total \\ 
 & & Non-MBA\\
\hline
 Promotional\\status & & \\ 
 Promoted & 12 & 48 & 60 \\
Not\ promoted &18 & 22 & 40 \\
Total & 30 & 70 & 100
\end{array}"

1. "P(MBA)=\dfrac{30}{100}=0.3"


2. "P(promoted)=\dfrac{60}{100}=0.6"


3. "P(promoted|MBA)=\dfrac{12}{30}=0.4"


4. "P(MBA|promoted)=\dfrac{12}{60}=0.2"


5. "P(promoted\cap MBA)=\dfrac{12}{100}=0.12"